# Deconvoluting Thermomechanical Effects in X-ray Diffraction Data using Machine Learning

### v1

@@ 3.4 X-ray Diffraction Simulations
The forward modeling framework for generating synthetic diffraction patterns for model training utilizes HEXRD [37, 38] and is described in more detail by [39]. Simulations build on our previous effort [20], but have now been modified to include the anisotropic scattering effects of elastic strains on the diffraction simulations. The nine elastic strain $\bm{\varepsilon}^{E}(\bm{r})$ and thermal strain fields $\varepsilon^{T}(\bm{r})$ (remapped to a 20 $\mu$m grid) calculated from thermomechanical simulations were used as input for the diffraction simulations. A scattering volume is built around each field point in which discrete scattering crystals can be placed. Here grains are modeled with a 25 $\mu$m equivalent grains size so 2 crystals are inserted around each scattering volume. Each crystal is modeled with 1 of misorientation to provide a realistic peak width. The crystals are randomly oriented which is consistent with the lack of texture found in the heat-treated IN625 being modeled. An isostrain (Taylor) assumption is made in which the simulated macroscopic elastic and thermal strains can be projected onto the crystals.
To model the effects of thermal and elastic strains, the the reciprocal lattice vectors, $\bm{g}$, within each grain are stretch from a reference $\bm{g_{0}}$:
@@
The reference lattice parameter from which $\bm{g_{0}}$ were calculated was 3.5981 $\AA$. At each time step, diffraction from a set of lattice planes within a grain is determined to occur based on the incoming X-ray energy and bandwidth
@@
Diffraction simulations for each elasto-plasticity simulation were repeated three times with the beam placed at 20 $\mu$m, 40 $\mu$m, and 60 $\mu$m from the top of the specimen and the rastering laser beam. The purpose of this is two-fold. Performing simulations at different X-ray beam positions on the sample increased the amount of diffraction images generated for model training, and it helped to account for any uncertainty in beam placement with respect to the temperature field in the sample. In total, 27 sets of X-ray simulations and 75 diffraction images from each thermomechanical simulation were generated (2025 total) for GPR model training. Once the 2D diffraction patterns were simulated for the entire time series, each pattern was integrated azimuthally in the same fashion as the experimental data which is illustrated in Fig. 3. Note that the mismatch between synthetic and experimental data of the relative intensities of each peak within each color highlights the local variations of texture due to the combination of beam and grain size. After integration, background scattering was added to the same data of the same magnitude as that observed in the experiment.
@@ 4 Machine-Learning Model and Training
Gaussian Process Regression [40] was used for learning the mapping between diffraction data and the distributions of thermal and elastic strain present within the diffraction volume. The approach here is an extension of our previous effort to map isotropic diffraction ring evolution to temperature (thermal strain) fields present, to now map anisotropic diffraction ring evolution (both peak shifts and broadening) to elastic and thermal strain distributions present. GPR assumes a normal distribution of mapping functions, $f$, (here strain field metrics) from input data, $\bm{x}$, (here diffracted intensity distributions along different sample directions). The mean mapping function ($\bar{f}$) from the normal distribution is the prediction of the model. A natural benefit of the approach is that uncertainty is estimated from the variance of the mapping functions. The mapping functions learned are linear combinations of input training data $\bm{x}^{*}$, where the coefficients are dependent on the similarity of the training and input data. If input data are similar (here determined by Euclidean distance) to training data, that training data are more heavily weighted as dictated by the chosen covariance function (see below). In addition, if the input data for the model are not similar to any training data, the uncertainty of the model prediction increases.
@@
where $\bm{x_{a}}$ and $\bm{x_{b}}$ are two input data points, while $\alpha$ and $L$ control the decay rate for weights. As $\alpha$ decreases, training data of increased dissimilarity from the input data are incorporated into model predictions, while $L$ provides a further control if necessary. Here we use $L=1$ and $\alpha=1$. For GPR model training and testing, 26 of the data sets were used, while one (No. 13, see Table 2) was reserved for testing. We trained 16 different GPR models to learn mapping between the mean, maximum, minimum, and standard deviations (STDs) of $\varepsilon^{T}$, $\varepsilon^{E}_{xx}$, $\varepsilon^{E}_{xy}$, and $\varepsilon^{E}_{yy}$ and the diffraction data. For these simulations (and the experiment), $\varepsilon^{E}_{xy}$ is minimal across the sample, however, this still provides a further test of the model. In this diffraction geometry, the projections of $\bm{\varepsilon}^{E}$ along $\bm{g}$ are minimal and as such $\varepsilon^{E}_{zz}$, $\varepsilon^{E}_{xz}$, and $\varepsilon^{E}_{yz}$ are not probed. We note that the trained GPR models are not fitting analytic functions to the peaks, and thus not restricted to regularly shaped diffraction peaks (e.g., Gaussian or Lorentzian) and are capable of mapping ‘split’ peaks or those with long tails to underlying stain distributions [20].
@@
To examine the examine the accuracy of the predictions of the trained GPR models, the reserved testing data set was used as input for the trained models and the output predictions were then compared to the true thermomechanical distributions quantities. Fig. 8 shows the comparisons between mean strains in diffraction volumes from the thermomechanical model output (here serving as ground truth) compared to predictions from the trained GPR model using diffraction data as input. Figs. 8a, 8b, 8c. 8d correspond to $\varepsilon^{T}$, $\varepsilon^{E}_{xx}$, $\varepsilon^{E}_{yy}$, and $\varepsilon^{E}_{xy}$. Note the difference in strain scales between thermal and elastic strains (due to large differences in magnitude) which will continue through the rest of the work. In general, the GPR models predictions for mean strains are accurate with the ground truth falling with the uncertainty bounds. In Fig. 8a, the thermal strain $\varepsilon^{T}_{Mean}$ predictions tend to be closer to the ground truth at room temperature, with deviations at higher strains (temperatures). For the primary mechanical strains $\varepsilon^{E}_{xx-Mean}$ in 8b, the opposite is true, the accuracy tends to be better at larger strains. The predictions of the mean $\varepsilon^{E}_{yy-Mean}$ and $\varepsilon^{E}_{xy-Mean}$ generally fit well withing the strains bounds where there is data (sans a small number of outliers), but we note that the magnitude of mean strains is relatively low for these strain components.
@@
Figure 8: Accuracy predicting reserved testing data of trained surrogate models to determine the mean of a) $\varepsilon^{T}$, b) $\varepsilon^{E}_{xx}$, c) $\varepsilon^{E}_{yy}$, and d) $\varepsilon^{E}_{xy}$ within a diffraction volume from input diffraction spectrum. The dashed blue line denotes one-to-one correlation, and the black line is a linear fit.
@@
Often of more interest are the maximum and minimum strains, particularly the maximum, within the diffraction volume. Figure 9 shows comparisons between maximum strains in the reserved testing data and the various GPR model predictions. For $\varepsilon^{T}_{Max.}$ in Fig. 9a, there is a small under prediction of the maximum strain at higher strain (temperature values). This is similar to $\varepsilon^{E}_{xx-Max.}$ in Fig. 9b which also show some degree of under prediction at higher strain values. In Figs. 9c and 9d, it can be seen that there is a small under-prediction of $\varepsilon^{E}_{yy-Max.}$ and $\varepsilon^{E}_{xy-Max.}$ respectively across the diffraction ranges of strain shown. The under predictions across all strain components may be related to very small volume fractions of material in general contributing the diffraction peaks in comparison to the mean. Similarly to Fig. 9, Fig. 10 shows comparisons between minimum strains in the reserved testing data and the various GPR model predictions. The predictions of the minimum thermal strains $\varepsilon^{T}_{Min.}$ in Fig. 10a generally match well to the ground truth across the total strain range. For the minimum $\varepsilon^{E}_{xx}$ strains, the trends differ from predictions of the maximum in that there is a small over prediction across the strain range. For $\varepsilon^{E}_{yy-Min.}$ and $\varepsilon^{E}_{xy-Min.}$, the strains are generally slightly under-predicted at high strain values and over predicted at lower strain values as seen Figs. 10c and 10d.
@@
Figure 9: Accuracy predicting reserved testing data of trained surrogate models to determine the maximum of a) $\varepsilon^{T}$, b) $\varepsilon^{E}_{xx}$, c) $\varepsilon^{E}_{yy}$, and d) $\varepsilon^{E}_{xy}$ within a diffraction volume from input diffraction spectrum. The dashed blue line denotes one-to-one correlation, and the black line is a linear fit.
@@
Figure 10: Accuracy predicting reserved testing data of trained surrogate models to determine the minimum of a) $\varepsilon^{T}$, b) $\varepsilon^{E}_{xx}$, c) $\varepsilon^{E}_{yy}$, and d) $\varepsilon^{E}_{xy}$ within a diffraction volume from input diffraction spectrum. The dashed blue line denotes one-to-one correlation, and the black line is a linear fit.
@@
The final strain metric for which GPR models were trained to extract from the diffraction data was the standard deviation (STD) of the distributions of strain within a diffraction volume. The results of the trained models for STD for $\varepsilon^{T}$, $\varepsilon^{E}_{xx}$, $\varepsilon^{E}_{yy}$, and $\varepsilon^{E}_{xy}$ are shown in Figs. 11a, 11b, 11c, and 11d respectively. The predictions of STD across GPR models across all strain types and components tend to under predict the spread of the temperature distributions. This again is likely due to the very small contributions of extreme values (particularly maximums) which contribute very little intensity to the diffraction pattern. However, the GPR predictions and In addition, the GPR models are not predicting any aphysical standard deviation values (such as a negative value).
@@
Figure 11: Accuracy predicting reserved testing data of trained surrogate models to determine the standard deviation of a) $\varepsilon^{T}$, b) $\varepsilon^{E}_{xx}$, c) $\varepsilon^{E}_{yy}$, and d) $\varepsilon^{E}_{xy}$ within a diffraction volume from input diffraction spectrum. The dashed blue line denotes one-to-one correlation, and the black line is a linear fit.
@@ 5 Application to Experimental Data
The trained GPR surrogate models were used to analyze the evolution of the thermal and elastic strains within the diffraction volume during the in situ laser melting experiment described in §2. As mentioned, the raw experimental diffraction images were azimuthally binned into four different 1D line profiles that were concatenated into a single vector (Fig 3). Results from applying the model to a single laser pass are presented, but applying the trained surrogate models to a second laser pass at a different position on the same sample are provided in an Appendix. Figure 12 shows the output metrics associated with thermal strains in the diffraction volume during the in situ experiment. The mean of the GPR predictions from each diffraction measurement are shown with black dots, while the red error bars are the variance of the GPR predictions which can be used as a measure of uncertainty. The mean (Fig. 12a), maximum (Fig. 12b), and minimum (Fig. 12c) of distribution of thermal strains all show a spike in thermal strain as the laser passes over the diffraction volume at 0.2 s. In addition, the standard deviations of the distributions (Fig. 12d) also rapidly increase as the laser passes over the specimen and then decreases back to near zero which is expected as the entire specimen returns back to room temperature.
@@
Figure 13 shows the output distribution metrics from the trained GPR models for $\varepsilon^{E}_{xx}$. In the $\varepsilon^{E}_{xx}$ output metrics, there appears to be a small non-zero tensile strain at the beginning of the test (possibly from sample mounting or manufacture) which then becomes compressive as the laser passes over the sample. As the sample cools, the expected tensile residual stress develops with a larger magnitude than the original tensile strain at the beginning of the test. The general behavior of these strain distribution metrics output from the GPR model is consistent with the evolution of strains expected from simulation (see Fig. 7). The most critical observation is that the GPR surrogate models appear to be able to isolate the relatively small (in comparison to the thermal strains) compressive mechanical strains that occur due to localized heating at the same time as capturing the large thermal strains (Fig. 12a). In addition, the standard deviation of $\varepsilon^{E}_{xx}$ shown in Fig. 12d increased from the start of the test and did not decay as the temperature fell, reflecting the distributions of residual strain in the specimens.
@@
The GPR model distribution metric outputs from the other two strain components, $\varepsilon^{E}_{yy}$ and $\varepsilon^{E}_{xy}$, are shown in Fig. 14 and Fig. 15. In the distributions in Fig. 14, a divergence of behavior is observed between the maximum (Fig. 14b) and minimum strains (Fig. 14c) in the distribution. While the maximum strains become slightly tensile as the laser passes over the specimen, the minimum strains become negative and then remain negative through cooling. The final mean of the distribution of $\varepsilon^{E}_{yy}$ is consistent with unconstrained Poisson contraction from the tensile residual strains that developed. Similar to $\varepsilon^{E}_{xx-STD}$, $\varepsilon^{E}_{yy-STD}$ peaks as the laser passes over the sample then decays to a value marginally larger than at the start of the test (Fig. 15d). The distributions metrics of $\varepsilon^{E}_{xy}$ generally show similar trends to that of $\varepsilon^{E}_{yy}$ as can be see in Fig. 15. However, the standard deviation of the distribution of shear strains $\varepsilon^{E}_{xy-STD}$ is at a minimum as the laser passes over the diffraction volume.
Figure 14: The evolving a) mean $\varepsilon^{E}_{yy-Mean}$ b) maximum $\varepsilon^{E}_{yy-Max.}$, c) minimum $\varepsilon^{E}_{yy-Min.}$, and d) standard deviation $\varepsilon^{E}_{yy-STD}$ of the distribution of elastic strain within the experimental X-ray diffraction volume with respect to time $t$ extracted using trained GPR surrogate models. The red error bars correspond to the square root of the variance (standard deviation) of the GPR surrogate model predictions.
Figure 15: The evolving a) mean $\varepsilon^{E}_{xy-Mean}$ b) maximum $\varepsilon^{E}_{xy-Max.}$, c) minimum $\varepsilon^{E}_{xy-Min.}$, and d) standard deviation $\varepsilon^{E}_{xy-STD}$ of the distribution of elastic strain within the experimental X-ray diffraction volume with respect to time $t$ extracted using trained GPR surrogate models. The red error bars correspond to the square root of the variance (standard deviation) of the GPR surrogate model predictions.

### v3

@@ 3.3 X-ray Diffraction Simulations
The forward modeling framework for generating synthetic diffraction patterns for model training utilizes HEXRD [29, 30] and is described in more detail by [31]. Simulations build on our previous effort [20], but have now been modified to include the anisotropic scattering effects of the deviatoric portion of the elastic strain tensor on the diffraction simulations. The series of nine elastic strain $\bm{\varepsilon}^{E}(\bm{r})$ and thermal strain fields $\varepsilon^{T}(\bm{r})$ (remapped to a 20 $\mu$m grid) calculated from thermomechanical simulations were used as input for the diffraction simulations. A scattering volume is built around each field point in which discrete scattering crystals can be placed. Here, grains are modeled with a 25 $\mu$m equivalent grain size so 2 crystals are inserted around each scattering volume for over 200 crystals in the diffraction volume. Each crystal is modeled with 1 of misorientation to provide a finite peak width perpendicular to the radial direction on the detector and ease numeric issues associated with calculating the diffraction condition. The crystals are randomly oriented which is consistent with the lack of texture found in the heat-treated IN625 being modeled. Here we project the macroscopic elastic and thermal strains onto the discrete diffracting crystals.
To model the effects of thermal and elastic strains, the reciprocal lattice vectors, $\bm{g}$, within each grain are stretched from a reference $\bm{g_{0}}$:
@@
The reference lattice parameter used for calculation of $\bm{g_{0}}$ was 3.5981 $\AA$. At each time step, diffraction from a set of lattice planes within a grain is determined to occur based on the incoming X-ray energy and bandwidth
@@
Diffraction simulations for each elasto-plasticity simulation were repeated three times with the beam placed at 20 $\mu$m, 40 $\mu$m, and 60 $\mu$m from the top of the specimen and the rastering laser beam (see Fig. 2). Note that the experiment was only performed with the beam centered at 20 $\mu$m from the sample surface. The purpose of this is two-fold. Performing simulations at different X-ray beam positions on the sample increased the amount of diffraction images generated for model training, and it helped to account for any uncertainty in beam placement with respect to the temperature field in the sample. In total, 27 sets of X-ray simulations in conditions similar to the experiment (3 different diffraction volumes for each of the 9 laser conditions simulated) were performed, with 75 diffraction images from each. This produced 2025 diffraction images total for GPR model training. Once the 2D diffraction patterns were simulated for the entire time series, each pattern was integrated azimuthally in the same fashion as the experimental data which is illustrated in Fig. 4. Note that the mismatch between synthetic and experimental data of the relative intensities of each peak within each color highlights the local variations of texture due to the combination of beam and grain size. After integration, background scattering was added to the same data of the same magnitude as that observed in the experiment.
@@ 4 Machine-Learning Model and Training
Gaussian Process Regression [32] was used for learning the mapping between diffraction data and the distributions of thermal and elastic strain present within the diffraction volume. The approach here is an extension of our previous effort of mapping isotropic diffraction ring evolution to temperature (thermal strain) fields present, to now map anisotropic diffraction ring evolution (both peak shifts and broadening) to elastic and thermal strain distributions present. GPR assumes a normal distribution of mapping functions, $f$, (here strain field metrics) from input data, $\bm{x}$, (here diffracted intensity distributions along different sample directions). The mean mapping function ($\bar{f}$) from the normal distribution is the prediction of the model. A natural benefit of the approach is that uncertainty is estimated from the variance of the mapping functions. Other ML approaches, such as neural networks, do not as readily provide measures of the uncertainty of predictions. The mapping functions learned are linear combinations of input training data $\bm{x}^{*}$, where the coefficients are dependent on the similarity of the training and input data. If input data are similar (here determined by Euclidean distance) to training data, those training data are more heavily weighted as dictated by the chosen covariance function (see below). In addition, if the input data for the model are not similar to any training data, the uncertainty of the model prediction increases.
@@
where $\bm{x_{a}}$ and $\bm{x_{b}}$ are two input data points, while $\alpha$ and $L$ control the decay rate for weights. As $\alpha$ decreases, training data of increased dissimilarity from the input data are incorporated into model predictions, while $L$ provides a further control if necessary. Here, we use $L=1$ and $\alpha=1$. The choice of rational quadratic kernel slows the decay of the interpolation leading to more training data points being used in each prediction.
@@
For GPR model training and testing, 26 of the data sets were used, while one (No. 13, see Table 2) was reserved for testing. We trained 16 different GPR models to learn mapping between the mean, maximum, minimum, and standard deviations (STDs) of $\varepsilon^{T}$, $\varepsilon^{E}_{xx}$, $\varepsilon^{E}_{xz}$, and $\varepsilon^{E}_{zz}$ within the illuminated diffraction volume and the diffraction data. For these simulations (and the experiment), $\varepsilon^{E}_{xz}$ is minimal across the sample, however, this still provides a further test of the model. In this diffraction geometry, the projections of $\bm{g}$ along $\bm{y}$ are minimal and as such $\varepsilon^{E}_{yy}$, $\varepsilon^{E}_{xy}$, and $\varepsilon^{E}_{yz}$ are not probed. We note that the trained GPR models are not fitting analytic functions to the peaks, and thus not restricted to regularly shaped diffraction peaks (e.g., Gaussian or Lorentzian) and are capable of mapping ‘split’ peaks or those with long tails to underlying strain distributions [20].
@@
To examine the accuracy of the predictions of the trained GPR models, the reserved testing data set was used as input for the trained models and the output predictions were then compared to the true thermomechanical distributions quantities. Fig. 7 shows the comparisons between mean strains in diffraction volumes from the thermomechanical model output (here serving as ground truth) compared to predictions from the trained GPR model using diffraction data as input. Figs. 7a, 7b, 7c, and 7d correspond to $\varepsilon^{T}$, $\varepsilon^{E}_{xx}$, $\varepsilon^{E}_{zz}$, and $\varepsilon^{E}_{xz}$. Note the difference in strain scales between thermal and elastic strains (due to large differences in magnitude) which will continue through the rest of the work. In general, the GPR models predictions for mean strains are accurate with the ground truth falling within the uncertainty bounds. In Fig. 7a, the thermal strain $\varepsilon^{T}_{Mean}$ predictions tend to be closer to the ground truth at room temperature, with deviations at higher strains (temperatures). For the primary mechanical strains $\varepsilon^{E}_{xx-Mean}$ in 7b, the opposite is true, the accuracy tends to be better at larger strains. The predictions of the mean $\varepsilon^{E}_{zz-Mean}$ and $\varepsilon^{E}_{xz-Mean}$ generally fit well within the strains bounds where there is data (besides a small number of outliers), but we note that the magnitude of mean strains is relatively low for these strain components.
@@
Figure 7: Accuracy predicting reserved testing data of trained surrogate models to determine the mean of a) $\varepsilon^{T}$, b) $\varepsilon^{E}_{xx}$, c) $\varepsilon^{E}_{zz}$, and d) $\varepsilon^{E}_{xz}$ within a diffraction volume from input diffraction spectra. The dashed blue line denotes one-to-one correlation, and the black line is a linear fit.
@@
Often of more interest are the maximum and minimum strains, particularly the maximum, within the diffraction volume. Figure 8 shows comparisons between maximum strains in the reserved testing data and the various GPR model predictions. For $\varepsilon^{T}_{Max.}$ in Fig. 8a, there is a small under-prediction of the maximum strain at higher strain (temperature values). This is similar to $\varepsilon^{E}_{xx-Max.}$ in Fig. 8b which also show some degree of under-prediction at higher strain values. In Figs. 8c and 8d, it can be seen that there is a small under-prediction of $\varepsilon^{E}_{zz-Max.}$ and $\varepsilon^{E}_{xz-Max.}$ respectively across the diffraction ranges of strain shown. The under-predictions across all strain components may be related to very small volume fractions of material in general contributing the diffraction peaks in comparison to the mean. Similarly to Fig. 8, Fig. 9 shows comparisons between minimum strains in the reserved testing data and the various GPR model predictions. The predictions of the minimum thermal strains $\varepsilon^{T}_{Min.}$ in Fig. 9a generally match well to the ground truth across the total strain range. For the minimum $\varepsilon^{E}_{xx}$ strains, the trends differ from predictions of the maximum in that there is a small over-prediction across the strain range. For $\varepsilon^{E}_{zz-Min.}$ and $\varepsilon^{E}_{xz-Min.}$, the strains are generally slightly under-predicted at high strain values and over-predicted at lower strain values as seen Figs. 9c and 9d.
@@
Figure 8: Accuracy predicting reserved testing data of trained surrogate models to determine the maximum of a) $\varepsilon^{T}$, b) $\varepsilon^{E}_{xx}$, c) $\varepsilon^{E}_{zz}$, and d) $\varepsilon^{E}_{xz}$ within a diffraction volume from input diffraction spectra. The dashed blue line denotes one-to-one correlation, and the black line is a linear fit.
@@
Figure 9: Accuracy predicting reserved testing data of trained surrogate models to determine the minimum of a) $\varepsilon^{T}$, b) $\varepsilon^{E}_{xx}$, c) $\varepsilon^{E}_{zz}$, and d) $\varepsilon^{E}_{xz}$ within a diffraction volume from input diffraction spectra. The dashed blue line denotes one-to-one correlation, and the black line is a linear fit.
@@
The final strain metric for which GPR models were trained to extract from the diffraction data was the standard deviation (STD) of the distributions of strain within a diffraction volume. The results of the trained models for STD for $\varepsilon^{T}$, $\varepsilon^{E}_{xx}$, $\varepsilon^{E}_{zz}$, and $\varepsilon^{E}_{xz}$ are shown in Figs. 10a, 10b, 10c, and 10d respectively. The predictions of STD across GPR models across all strain types and components tend to under-predict the spread of the temperature distributions. This again is likely due to the very small contributions of extreme values (particularly maximums) which contribute very little intensity to the diffraction pattern. However, the GPR models are not predicting any aphysical standard deviation values (such as a negative value).
@@
Figure 10: Accuracy predicting reserved testing data of trained surrogate models to determine the standard deviation of a) $\varepsilon^{T}$, b) $\varepsilon^{E}_{xx}$, c) $\varepsilon^{E}_{zz}$, and d) $\varepsilon^{E}_{xz}$ within a diffraction volume from input diffraction spectra. The dashed blue line denotes one-to-one correlation, and the black line is a linear fit.
@@ 5 Application to Experimental Data
The trained GPR surrogate models were used to analyze the evolution of the thermal and elastic strains within the diffraction volume during the in situ laser melting experiment described in §2. As mentioned, the raw experimental diffraction images were azimuthally binned into six different 1D line profiles that were concatenated into a single vector (Fig. 4). Results from applying the model to a single laser pass are presented, but applying the trained surrogate models to a second laser pass at a different position on the same sample are provided in Appendix A2. Figure 11 shows the output metrics associated with thermal strains in the diffraction volume during the in situ experiment. The mean of the GPR predictions from each diffraction measurement are shown with black dots, while the red error bars are the variance of the GPR predictions which can be used as a measure of uncertainty. The mean (Fig. 11a), maximum (Fig. 11b), and minimum (Fig. 11c) of distribution of thermal strains all show a spike in thermal strain as the laser passes over the diffraction volume at 0.2 s. In addition, the standard deviations of the distributions (Fig. 11d) also rapidly increase as the laser passes over the specimen and then decreases back to near zero which is expected as the entire specimen returns back to room temperature.
@@
Figure 12 shows the output distribution metrics from the trained GPR models for $\varepsilon^{E}_{xx}$. In the $\varepsilon^{E}_{xx}$ output metrics, there appears to be a small non-zero tensile strain at the beginning of the test (possibly from sample mounting or manufacture) which then becomes compressive as the laser passes over the sample. As the sample cools, the expected tensile residual stress develops with a larger magnitude than the original tensile strain at the beginning of the test. The general behavior of these strain distribution metrics output from the GPR model is consistent with the evolution of strains expected from simulation (see Fig. 6). The most critical observation is that the GPR surrogate models appear to be able to isolate the relatively small (in comparison to the thermal strains) compressive mechanical strains that occur due to localized heating, while simultaneously capturing the large thermal strains (Fig. 11a). In addition, the standard deviation of $\varepsilon^{E}_{xx}$ shown in Fig. 11d increased from the start of the test and did not decay as the temperature fell, reflecting the distributions of residual strain in the specimens.
@@
The GPR model distribution metric outputs from the other two strain components, $\varepsilon^{E}_{zz}$ and $\varepsilon^{E}_{xz}$, are shown in Fig. 13 and Fig. 14. In the distributions in Fig. 13, a difference of behavior is observed between the maximum (Fig. 13b) and minimum strains (Fig. 13c) in the distribution. While the maximum strains become slightly tensile as the laser passes over the specimen, the minimum strains become negative and then remain negative through cooling. The final mean of the distribution of $\varepsilon^{E}_{zz}$ is consistent with unconstrained Poisson contraction from the tensile residual strains that developed. Similar to $\varepsilon^{E}_{xx-STD}$, $\varepsilon^{E}_{zz-STD}$ peaks as the laser passes over the sample then decays to a value marginally larger than at the start of the test (Fig. 14d). The distributions metrics of $\varepsilon^{E}_{xz}$ generally show similar trends to that of $\varepsilon^{E}_{zz}$ as can be see in Fig. 14. However, the standard deviation of the distribution of shear strains $\varepsilon^{E}_{xz-STD}$ is at a minimum as the laser passes over the diffraction volume.
Figure 13: The evolving a) mean $\varepsilon^{E}_{zz-Mean}$ b) maximum $\varepsilon^{E}_{zz-Max.}$, c) minimum $\varepsilon^{E}_{zz-Min.}$, and d) standard deviation $\varepsilon^{E}_{zz-STD}$ of the distribution of elastic strain within the experimental X-ray diffraction volume with respect to time $t$ extracted using trained GPR surrogate models. The red error bars correspond to the square root of the variance (standard deviation) of the GPR surrogate model predictions.
Figure 14: The evolving a) mean $\varepsilon^{E}_{xz-Mean}$ b) maximum $\varepsilon^{E}_{xz-Max.}$, c) minimum $\varepsilon^{E}_{xz-Min.}$, and d) standard deviation $\varepsilon^{E}_{xz-STD}$ of the distribution of elastic strain within the experimental X-ray diffraction volume with respect to time $t$ extracted using trained GPR surrogate models. The red error bars correspond to the square root of the variance (standard deviation) of the GPR surrogate model predictions.